Circular No. 9094

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

η CARINAE

K. Davidson and A. Mehner, University of Minnesota; and J. C. Martin, University of Illinois at Springfield, on behalf of a group using the Hubble Space Telescope, report that the central star in η Car has recently brightened to V = 5.1. This includes a region only 0′′.3 across, in spectra obtained with the Space Telescope Imaging Spectrograph (STIS) on 2009 Aug. 19. The star now accounts for half of the total light seen in the Homunculus nebula, compared to less than 10 percent before 1995. In view of the central star’s 0.6-magnitude brightening since its 2006 peak (Martin et al. 2006, A.J. 132, 2717), unsteadiness in 2007–2008, and 0.5-magnitude increase in mid-2009, η Car currently merits frequent attention by additional

observers. See also http://etacar.fcaglp.unlp.edu.ar/.

SATELLITE OF JUPITER Further to IAUC 8826, J. S. Blue reports that the IAU Working Group

for Planetary System Nomenclature has approved a new designation and name for the 50th confirmed satellite of Jupiter — that for S/2003 J 17, which was re-discovered this past August (cf. MPEC 2009-S76):

Jupiter L Herse = S/2003 J 17  IAUC 8116

(50000) QUAOAR I The IAU Committee on Small Bodies Nomenclature has approved the

name “Weywot” for the satellite of the transneptunian minor planet (50000) Quaoar that was announced on IAUC 8812 (cf. MPC 67220).

COMET 228P/LINEAR The permanent number 228P has been assigned to comet C/2009 U2

= P/2001 YX_{127} (cf. IAUC 9085; MPC 67349).

COMET C/2009 M3 (SOHO) Further to IAUC 9092, another Kreutz sungrazer has been found on

SOHO website images. C/2009 M3 was stellar in appearance (peak mag ∼ 6.5–7) on C3 images, but slightly diffuse and elongated in C2 images.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 M3 June 22.700 5^{h}57.^{m}8 +20^{o}35′ C3/2 BZ 2009-P02
2009 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green